# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

### v5

@@ 5.5 Performance on Computer Vision
Cheon [20] validated the effectiveness of KAN in visual tasks on the MNIST, CIFAR10, and CIFAR100 datasets. The result shows that KAN has a better performance than MLP on CIFAR10, and CIFAR100 but slightly worse than the state-of-the-art ResNet-18. Bodner et al [20] create Convolutional Kolmogorov-Arnold Networks (Convolutional KANs) and validate the performance of Convolutional KANs against traditional architectures across Fashion-MNIST dataset. Their experiments show that Convolutional KANs reach the a similar level of accuracy with half parameters which demonstrate the potential of KANs in computer version field. However, Tran et al [21] explore the limits of KAN on classification tasks. They claim that KAN exceed MLP despite theoretical advantages in accuracy and interpretability since KAN will consume more resources and have higher latency than MLPs. Their results show that MLP is still the better option when it comes to classification task.
@@ 6.2 Interpretability and Symbolic Regression
Monotonic Kolmogorov-Arnold Network (MonoKAN)
@@
Bayesian Kolmogorov Arnold Networks (BKANs)
@@
7 Health and Medical Field
@@
8 Computer Vision and Image Processing
@@
Residual Kolmogorov-Arnold Network (RKAN)
@@
9 Graph Neural Network and Collaborative Filtering
@@
10 Physics-Informed Neural Networks and Engineering Applications
@@ 11.2 Other Application and Study of KANs
Geographic Information Science
@@
Yu et al [54] explore the KANs for image quality assessment by comparing KANs with different activation functions. After rigorous and detailed comparative experiments, they introduce TaylorKAN with mid-level feature input, which shows great performance on 3 realistic image databases (BID2011, CID2013 and CLIVE).

### v6

@@ 5.5 Performance on Computer Vision
Cheon [20] validated the effectiveness of KAN in visual tasks on the MNIST, CIFAR10, and CIFAR100 datasets. The result shows that KAN has a better performance than MLP on CIFAR10, and CIFAR100 but slightly worse than the state-of-the-art ResNet-18. Bodner et al [20] create Convolutional Kolmogorov-Arnold Networks (Convolutional KANs) and validate the performance of Convolutional KANs against traditional architectures across Fashion-MNIST dataset. Their experiments show that Convolutional KANs reach the a similar level of accuracy with half parameters which demonstrate the potential of KANs in computer version field.Han et al [21] have proved that KAN has potential in face recognition field compared with KAN and they propose 2 different variants. However, Tran et al [22] explore the limits of KAN on classification tasks. They claim that KAN exceed MLP despite theoretical advantages in accuracy and interpretability since KAN will consume more resources and have higher latency than MLPs. Their results show that MLP is still the better option when it comes to classification task.Moreover, Mohan et al [23] highlight that KAN is sensitive to hyperparameters and extremely demanding on computing resources and it need justify its structure to fit large computer vision problems.
@@ 6.2 Interpretability and Symbolic Regression
MonoKAN (Monotonic Kolmogorov-Arnold Network)
@@
BKANs (Bayesian Kolmogorov Arnold Networks )
@@
6.3 Health and Medical Field
@@
ECG-SleepNet
Aghaomidi and Wang present a normal approach to classify the different sleep stages. [41] They use KAN to distinguish the five sleep stages after collecting different biological signals. The best accuracy is about 81 percent, which shows great advance compared with previous models.
MvKeTR (Multi-view perception Knowledge-enhanced TansfoRmer)
Deng et al [42] designed MvKeTR, a new framework which combines Multi-View Per-ception Aggregator (MVPA), Transformer and KAN to help generate CT report. In ablation experiment, they show that MvKeTR has great enhancement compared with previous frameworks.
KAN-Mamba FusionNet
Agrawal et al [43] combine KAN with Mamba layers for medical image segmentation. In their experiments, they compared their framework on BUSI, Kvasir-Seg and GlaS datasets. They claim that their model outperforms the current state-of-the-art in both IoU and F1 scores.
6.4 Computer Vision and Image Processing
@@
RKAN (Residual Kolmogorov-Arnold Network)
@@
6.5 Graph Neural Network and Collaborative Filtering
@@
6.6 Physics-Informed Neural Networks and Engineering Applications
@@ 7.2 Other Application and Study of KANs
7.3 Geographic Information Science
@@
Yu et al [59] explore the KANs for image quality assessment by comparing KANs with different activation functions. After rigorous and detailed comparative experiments, they introduce TaylorKAN with mid-level feature input, which shows great performance on 3 realistic image databases (BID2011, CID2013 and CLIVE). Cang et al [60] compared KAN with its variant, Convolutional KAN (CKAN), which are popular in this field. However, they claim that CKAN underperforms traditional CNNs due to low generalization, while KAN, despite strong fitting, also lacks generalization, making it unsuitable for replacing the final layer with an NLP layer.
Medical Field
@@
Neuroanatomy
In order to solve the problem brought by 3-hinge gyrus identification, Chen et al [62] introduce a new framework built with KANs. In their experiments, they show the strong ability in typology.They claim that the framework can play an important role in neuroanatomy field.
@@ 7.7 Improvements and Optimizations
PowerKAN
In order to solve the problems caused by the high time complexity of KAN, Qiu et al [73] introduce PowerKAN. They claim PowerKAN accelerates the speed of computing in approximately the same training time as MLP and shows stronger expressive power than KAN.
Constraint Informed Kolmogorov-Arnold Networks (CIKAN)
Kim et al [74] combine Constrained-Informed Neural Network (CINN) with KAN and propose CIKAN. They validate CIKAN-TSG through simulations of constrained spacecraft rendezvous on elliptic orbits, comparing it with MLP-CINNs and standard TSGs.